OpenStudy (anonymous):

Use addition of algebraic fractions to simplify the given expression. (1/x − 2)-(8/x^2+4x-12)

OpenStudy (anonymous):

\[\left(\frac{1}{x}-2\right)-\left(\frac{8}{x^2}+4x-12\right)\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

or \[\left(\frac{1}{x-2}\right)-\left(\frac{8}{x^2+4x-12}\right)\]

OpenStudy (anonymous):

Just factor the bottom of the second fraction and you will see that you will get (x-2)(x+6). This is your least common denominator. Multiply everything by that, and you should end up with 1(x+6) + 8